What is the length of the hypotenuse of the right triangle ABC if AC=6 and AD=5?
Answer by RAY100(1637)

The hypotenuse of a right triangle is the longest side.
.
Normally it is c,, while the shorter legs are, a & b
.
The Pythagorean Theorem lets, c^2 = a^2 + b^2
.
ASSUMING, a=5,,,and b=6,,,,,,(problem is confusing nomenclature)
.
c^2 = (5)^2 +(6)^2
.
c^2 = 25 +36 = 61,,,,,,,taking square root of both sides
.
c= sqrt(61) = 7.81
.
.
check with calc,,,7.81^2 = 5^2 +6^2 = 61,,,,ok
.